four children had three boxes of coloured pencils. They decided to open all the three boxes, to shair the pencils fairly. There were 52 pencils in each box. How many pencils did each child get ? Now look at the attempt from grade 4 learner to solve this problem .

a ) briefly describe how the learners solved this problem.

b )use any strategy to judge the reasonableness of thr learner's solution.

1
Expert's answer
2021-07-08T06:48:41-0400

a) If we have four children and a box of 52 coloured pencils, we can divide them fairly "(\frac {52} {4} = 13)" or just manually give each child one pencil and so on.

When the first box is empty, each child will have 13 pencils.


b) We still have two boxes of pencils, but as they do have 52 pencils each as well, each child will get 13 pencils from the second box and 13 pencils from the third box.

To sum up, each child gets "13 + 13 + 13 = 39" pencils.

or

"=\frac {52×3}{4}=39" pencils.